OpenStudy (anonymous):

x^2 = 4(-2)y. how do i get this equation to equal "y"

OpenStudy (anonymous):

x^2 = -8y (x^2)/-8 = y

Directrix (directrix):

x^2 = 4(-2)y Is this x^2 = 4 ^ (-2y) or x^2 = 4 * (-2y)

OpenStudy (anonymous):

x^2 = -8y

Directrix (directrix):

x^2 = -8y -8y = x^2 (-1/8) (-8y) = (-1/8) (x^2) y = -(x^2) / 8

OpenStudy (anonymous):

thank you!